### v1

# Coherent coupling of YBCO superconducting resonators and 100-nm-thick YIG films

Abstract
In cavity magnonics, magnon-photon hybridization has been widely investigated for both fundamental studies and applications. Planar superconducting resonators operating at microwave frequencies have demonstrated the possibility to achieve high couplings with magnons by exploiting the confinement of the microwave field in a reduced volume. Here we report a study of the coupling of high-$T_{c}$ YBCO superconducting waveguides with 100-nm-thick YIG magnetic films. We study the evolution of mode frequencies at different temperatures and extract the coupling strength of hybrid magnon-photon modes. We show that the experimental results can be reproduced using a simple model in which the temperature dependence of the penetration depth accounts for the evolution of the polaritonic spectrum.
keywords:
Magnons , Spin waves , Microwaves , Strong coupling , Hybrid systems , Superconducting resonators
journal: Journal of Magnetism and Magnetic Materials
[…]
[label1]organization=Istituto Nanoscienze - CNR, Centro S3, addressline=via G. Campi 213/A, city=Modena, postcode=41125, country=Italy
[…]
[label2]organization=Dipartimento di Scienze Fisiche, Informatiche e Matematiche, Universita di Modena e Reggio Emilia, addressline=via G. Campi 213/A, city=Modena, postcode=41125, country=Italy
1 Introduction
Magnon-photon hybrids are obtained by coherently coupling collective excitations in coupled spin systems, that is spin waves and their quantized counterpart magnons, with microwave photon modes. These systems have proven to be fertile ground for both fundamental studies and applications. In the former case, topics like quantum magnonics [1, 2], nonlinear effects [3, 4, 5], ultrastrong coupling [6, 7, 8], non-Hermitian physics [9, 10], Floquet engineering [11], magnonic frequency combs [12] have been recently investigated. Applications in magnonics and quantum technologies, either at room temperature or in the quantum regime at mK temperatures, are related, but not restricted, to computation, sensing, microwave-to-optical transduction, coherent coupling of remote physical systems, emission of microwaves, among others [13, 14, 15, 16, 17].
For these studies, magnetic materials such as electrically insulating yttrium iron garnet (YIG) or conducting permalloy, which are characterized by high spin densities and low ferromagnetic resonance (FMR) damping rates, have found widespread attention [18]. The incorporation of these materials into superconducting circuits has opened additional possibilities [19, 20, 21, 22]. Microwave modes in planar superconducting resonators have been exploited to achieve strong [23, 24, 25, 26] and ultrastrong [27, 28, 29, 8, 30] coupling with magnons, also using nanomagnets fabricated directly onto the resonator [25, 26]. Furthermore, superconducting qubits were coupled to YIG spheres to detect single-magnon excitations [1, 31].
High critical temperature ($T_{c}$) superconducting resonators, in particular YBa2Cu3O7 (YBCO), show $T_{c}$ up to about 90 K and resilience in applied magnetic field [32]. Strong and ultrastrong coupling with spin systems have been achieved, respectively, using paramagnetic ensembles [33, 34, 35, 36] and 5-$\mu\mathrm{m}$-thick YIG films [8, 30]. In the latter case, the evolution of polaritonic modes at temperature below $T_{c}$ shows a progressive frequency shift of the hybrid magnon-photon mode, which has been attributed to the effect of Meissner currents and correlated with the penetration depth in the superconducting layer [30].
Figure 1: (a) Sketch of the experimental geometry. (b) Spectral map showing the numerical derivative of broadband transmission spectra acquired at $T=30~{}\mathrm{K}$. Red and green dashed lines show $\omega_{0}$ and $\omega_{1}$ as derived respectively from Eq. 2 and Eq 4. The blue dashed line show $\omega_{FMR}$. (c) Evolution, as a function of the temperature, of the $S_{21}$ spectra measured at $\mu_{0}H_{0}=0.23~{}\mathrm{T}$. The incident microwave power is $-15~{}\mathrm{dBm}$. The background transmission, derived from high magnetic field data, was subtracted from the spectra. Solid lines display the Lorentzian fit (Eq. 1).
The spectrum of spin-wave excitations in the YIG film and the coupling with the resonator are dependent upon the thickness of the magnetic layer. We have experimentally investigated the interplay between YBCO and 100-$\mathrm{nm}$-thick YIG films. We first study the broadband spectrum of excitations acquired using a YBCO coplanar waveguide (CPW) transmission line at different temperatures. We then analyze the coupling between microwave and magnon modes using a CPW resonator. We finally discuss how the extracted physical quantities evolve by varying the thickness of the YIG film.
2 Broadband transmission spectroscopy
We studied a 100-nm-thick YIG film that was grown by liquid phase epitaxy on a 0.5-mm-thick Gadolinium Gallium Garnet (GGG) substrate. The magnetic film was positioned above the YBCO layer and gently pushed from the backside. Transmission ($S_{21}$) spectra were measured as a function of frequency using a broadband CPW line with the external magnetic field ($H_{0}$) applied parallel to the central conductor (Fig. 1(a)). At low temperature ($T=30~{}\mathrm{K}$), the transmission derivative ($\partial S_{21}/\partial H_{0}$) map evidences the presence of a main resonance mode at the lowest frequency ($\omega_{0}$) plus an additional mode ($\omega_{1}$), which has much weaker amplitude, shifted by approximately $1.2~{}\mathrm{GHz}$ with respect to the former (Fig. 1(b)).
As a function of frequency, transmission spectra taken at $\mu_{0}H_{0}=0.23~{}\mathrm{T}$ show well-defined absorption dips (Fig. 1(c)). As the temperature decreases, the resonance frequency, $\omega_{0}$, progressively moves toward higher frequencies. The spectral lineshape of the absorption dip, obtained by subtracting the background from the transmission spectra, can be fitted using a Lorentzian curve [37]
$\bar{S}_{21}=\frac{a}{1+\left(\frac{\omega-\omega_{0}}{\frac{1}{2}\Delta\omega}\right)^{2}},$ (1)
where $\omega$ is the frequency, $a$ is the amplitude and $\Delta\omega$ is the full width at half maximum linewidth (Fig. 1(c)). In particular, from the fits we note that $\Delta\omega$ increases from $30~{}\mathrm{MHz}$ at 80 K to $64~{}\mathrm{MHz}$ at 10 K. These values are influenced by the presence of both paramagnetic GGG substrate and superconducting YBCO layer as we discuss in the following.
Fig. 2 shows broadband spectral maps taken at temperatures between 80 and 10 K. In order to make the position of the YIG modes more evident, we plot the second derivative of the transmission with respect to the magnetic field $\partial^{2}S_{21}/\partial H_{0}^{2}$. We note that spectra taken above 50 K show the presence of $\omega_{0}$ only, while $\omega_{1}$ is visible at 30 K and below. $\omega_{1}(H_{0})$ curve varies only slightly from 10 K to 30 K (Fig. 2).
Figure 2: Evolution of broadband spectral maps measured at different temperatures. The color scale shows the second derivative of transmission with respect to the magnetic field. Dashed lines show the curves calculated with Eq. 2 and Eq. 4, as indicated.
Considering the lateral profile of the CPW broadband line and the Damon-Eshbach geometry in our experiment, we expect that the transmission line excites spin wave modes with wavenumber up to $k_{y}=3\times 10^{5}~{}\mathrm{rad~{}m}^{-1}\approx 2\pi/s$ [38, 30]. According to the analytical model by Kalinikos ans Slavin (KS) [39], the frequency of the lowest mode can be calculated as [40]
$\begin{split}&\omega_{0}/2\pi=\\ &=\mu_{0}\gamma\sqrt{H_{0}(H_{0}+M_{s})+(M_{s})^{2}P_{00}(k_{y}d)\left(1-P_{00}(k_{y}d)\right)},\end{split}$ (2)
where $\gamma=28.02~{}\mathrm{GHz/T}$ is the electron gyromagnetic ratio and $P_{00}=1+[(1-\exp(-k_{y}d))/k_{y}d]$. We note that, being $d=100~{}\mathrm{nm}$ and $k_{y}d=0.03$, Eq. 2 results very near to $\omega_{FMR}=\mu_{0}\gamma\sqrt{H_{0}(H_{0}+M_{s})}$ (Fig. 1(a)). The temperature dependence of Eq. 2 derives from the saturation magnetization of YIG, $M_{s}$, which follows [41, 42, 30]
$M_{s}=M_{0}(1-uT^{3/2}-vT^{5/2}),$ (3)
being $u=23\times 10^{-6}~{}\mathrm{K^{-3/2}}$ and $v=1.08\times 10^{-7}~{}\mathrm{K^{-5/2}}$ [42]. Using $\mu_{0}M_{0}=0.28$ T, we can reproduce the magnetic field dependence of $\omega_{0}$ in the entire temperature range (Fig. 1(a) and Fig. 2).
Higher spin wave resonances beyond $\omega_{0}$ are described in the KS model as Perpendicular Standing Spin Wave (PSSW) modes [39]. The frequency of the first PSSW mode approximately follows [40]
$\begin{split}&\omega_{1}/2\pi=\\ &=\gamma\mu_{0}\sqrt{\left[H_{0}+\frac{2A}{M_{s}}\left(k_{y}^{2}+\left(\frac{\pi}{d}\right)^{2}\right)\right]\left[H_{0}+\frac{2A}{M_{s}}\left(k_{y}^{2}+\left(\frac{\pi}{d}\right)^{2}\right)+M_{s}+H_{0}\left(\frac{Ms/H_{0}}{\pi/d}\right)^{2}k_{y}^{2}\right]},\end{split}$ (4)
where $A$ is the exchange constant of the YIG film. The experimental spectra in Figs. 1 and 2 can be reproduced by using $A=5.2~{}\mathrm{pJ/m}$ at 30 K and $A=5.5~{}\mathrm{pJ/m}$ at 10 K.
3 Coupling between magnetic film and resonator
The 100 nm thick YIG film was subsequently placed on top of the YBCO resonator to test the coupling between their modes. Transmission spectra were acquired between 10 K and 85 K with the film in contact with the resonator. Fig. 3 shows the formation of two polaritonic branches, indicating hybridization between the magnon and photon modes. At low temperature, the splitting of approximately 460 MHz corresponds to the collective coupling $g\approx 230~{}\mathrm{MHz}$. As the temperature increases up to the critical temperature of YBCO, panels (a-f) show a progressive lowering of the resonator frequency ($\omega_{c}$) along with a decrease in polariton splitting.
Figure 3: Transmission ($\partial S_{21}/\partial H_{0}$) spectra acquired at different temperatures. The YIG film is positioned in contact with the resonator and the incident microwave power is $-15~{}\mathrm{dBm}$. For comparison, spectra taken in the same conditions with the GGG face of the sample located above the resonator do not show any resonant mode. The calculated polaritonic modes $\Omega_{\pm}$ are represented by red dashed lines; blue and yellow dashed lines show $\omega_{0}$ and $\omega_{b}$, respectively.
In order to extract meaningful parameters from the experimental spectra and compare them with previous results, we follow the analysis reported in [30]. The upper ($\Omega_{+}$) and lower ($\Omega_{-}$) polaritons can be described as the effect of the coupling between the resonator and a single magnetic mode ($\omega_{b}$). Their evolution follows [8]
$\Omega_{\pm}=\frac{1}{\sqrt{2}}\sqrt{\omega_{c}^{2}+\omega_{b}^{2}\pm\sqrt{\left(\omega_{c}^{2}-\omega_{b}^{2}\right)^{2}+16\omega_{c}\omega_{b}g^{2}}}~{},$ (5)
where $\omega_{b}$ is correlated to both magnetic field and temperature assuming that
$\omega_{b}=\omega_{0}+\delta_{sc},$ (6)
where $\omega_{0}$ the frequency of the lowest YIG mode (Eq. 2) and $\delta_{sc}$ a temperature-dependent change. The latter can be quantified by self-consistently including the spin-wave induced Meissner currents in the Landau-Lifshitz-Gilbert equation, to obtain [21]
$\delta_{sc}\approx\gamma\mu_{0}M_{s}k_{y}dr\frac{1-e^{-\frac{2t}{\lambda_{L}}}}{(k_{y}\lambda_{L}+1)^{2}-(k_{y}\lambda_{L}-1)^{2}e^{-\frac{2t}{\lambda_{L}}}},$ (7)
where $t$ is the thickness of the YBCO film and $r$ is a dimensionless geometrical factor associated with the YIG thickness and spin-wave ellipticity. $\lambda_{L}$ is the penetration depth of YBCO, which in the simplest two-fluid binomial approximation reads [43]
$\lambda_{L}=\frac{\lambda_{L}(0)}{\sqrt{1-\left(\frac{T}{T_{c}}\right)^{p}}},$ (8)
where $T_{c}$ is the critical temperature, $\lambda_{L}(0)$, is the London penetration length in the zero temperature limit, and $p=4/3$ for d-wave superconductors [43, 44, 45].
Figure 4: Frequency of the resonator extracted from transmission maps acquired at different temperatures (circles). The solid line shows the fit with Eq. 9. Inset. Coupling strength and frequency shift calculated using, respectively, the resonator frequencies (circles) and Eq. 9 (solid lines).
By means of Eq. 5 we have fitted the evolution of polaritons in Fig. 3. The frequency of the resonator, $\omega_{c}(T)$, is used at each temperature as a free parameter; $\omega_{b}$ has been calculated using Eq. 7, where $r=0.45$ and the penetration depth has been derived from Eq. 8 with $T_{c}=86.1~{}\mathrm{K}$ and $\lambda_{L}(0)=97~{}\mathrm{nm}$. The latter parameters are obtained from the fit of $\omega_{c}(T)$ (Fig. 4) using [45]
$\omega_{c}\approx\omega_{c}(0)\sqrt{\frac{L[\lambda_{L}(T_{0})]}{L[\lambda_{L}(T)]}},$ (9)
being $L[\lambda_{L}(T)]$ the temperature-dependent inductance and $\omega_{c}(0)=9.7~{}\mathrm{GHz}$ at $T_{0}=10~{}\mathrm{K}$. Finally, the collective coupling $g=g_{s}\sqrt{2s_{\mathrm{Fe}}N_{s}}$, where $s_{\mathrm{Fe}}=5/2$ is the single-ion spin of Fe3+, is calculated assuming that the number of spins is $N_{s}=1.49\times 10^{13}$ in the entire temperature range. The temperature dependence of $g$ derives, through $\omega_{c}(T)$, from the spin-photon coupling
$g_{s}\approx\frac{\mu_{0}\omega_{c}}{4w}\sqrt{\frac{h}{Z_{0}}},$ (10)
where $b_{vac}$ is the vacuum magnetic field of the resonator and $Z_{0}=58~{}\mathrm{\Omega}$ is the nominal impedance of the CPW line [30].
In summary, Fig. 4 shows the temperature evolution of the parameters that describe the experimental data. From the fitted values of $\omega_{c}(T)$ and using the values of $\delta_{sc}(T)$ and $g(T)$ calculated using $r=0.45$ and $N_{s}=1.49\times 10^{13}$, we can reproduce the evolution of the polaritonic spectrum in the 10-85 K range (Fig. 3).
4 Discussion
Data acquired at low temperature using the YBCO broadband line (Fig. 1(b) and Fig. 2) evidence the presence of two magnetic modes having frequencies $\omega_{0}$ and $\omega_{1}$. Their evolution as a function of magnetic field and temperature has been analyzed using the KS model and taking into account the saturation magnetization ($M_{s}$) and exchange constant ($A$). The values found for such parameters (Sect. 2) are in line with those reported in the literature [42, 46]. However, because the magnetic film is in contact with the surface of the superconductor, additional effects emerge due to their interplay [30] as evidenced by the comparison of transmission spectra taken at different temperatures.
The evolution of broadband maps (Fig. 2) shows that $\omega_{1}$ becomes visible only at $T\leq 30~{}\mathrm{K}$. The saturation magnetization and exchange constant of YIG have a weak temperature dependence in this range; by contrast, the characteristics of the YBCO film are strongly temperature dependent. In particular, the penetration depth decreases at low $T$, where the microwave field becomes more efficiently confined to the surface of the transmission line. In these conditions, similar to [8, 30], we note that the transmission line becomes more sensitive to spin wave modes having a short decay distance of the stray field, such as for high wavenumbers [38, 47].
The linewidths extracted from the broadband spectra in Fig. 1(c) show values larger than what is typically reported for YIG/GGG films. Losses in YIG/GGG films have been widely investigated as a function of experimental parameters such as temperature, film thickness, and wavenumber [48, 49, 50]. At room temperature, FMR resonances show linewidths well below 1 MHz; at low temperature the broadening of spin wave modes has been attributed to effects like dipolar coupling to the partially magnetized GGG substrate and relaxation of rare earth ions, resulting in linewidths between 1 and 30 MHz. Moreover, studies of magnetic losses in YIG films having conductive electrodes deposited on top have shown increased values of linewidth [51, 52, 53]. In our case, the wide linewidths in Fig. 1(c) can be attributed to the presence of the superconductor. We just mention that, due to Meissner screening and nonperfect parallelism of the film surface with respect to magnetic field, we expect field inhomogeneity at the superconductor-magnetic film interface.
Further evidence of the interplay between magnetic and superconducting layers emerges in transmission spectra acquired using the superconducting resonator. In the first place, we note that $\omega_{1}$ is not visible in Fig. 3, therefore, we can rule out the effects of higher modes observed with three-dimensional cavities [54]. Following the results obtained with 5-$\mathrm{\mu m}$-thick YIG films [30], in Sect. 3 we have assumed that $\omega_{b}$ is derived by the shift of the lowest YIG mode, $\omega_{0}$ (Eq. 6), as quantified by Eq. 7. In particular, the latter predicts that $\delta_{sc}$ is proportional to the thickness of the YIG film ($d$): the maximum value of $\delta_{sc}$ obtained in Ref. [30] for $d=5~{}\mathrm{\mu m}$ exceeds $1~{}\mathrm{GHz}$ whereas in the present case ($d=100~{}\mathrm{nm}$) $\delta_{sc}$ drops to $0.1~{}\mathrm{GHz}$ (Fig. 4). However, we notice that the geometrical factor $r$ used to fit the spectra is different in the two cases. These parameters well reproduce the evolution of the transmission data in Fig. 3. The absence of a noticeable frequency shift in the broadband spectra (Fig. 2) could be related to the relatively wide linewidth of $\omega_{0}$ (Fig. 1(c)).
It is interesting to compare the coupling strength values ($g$) obtained with YIG films having different thicknesses, though similar lateral dimensions, and obtained using YBCO resonators having the same geometry. The maximum coupling strength derived from the spectra in Fig. 3 is $0.23~{}\mathrm{GHz}$ while the coupling obtained with a 5-$\mathrm{\mu m}$-thick YIG film is around $1.1~{}\mathrm{GHz}$ [30]. We first note that the ratio between the thicknesses is 50 while the coupling strength achieved with the thicker film is 5 times higher; considering that $g\propto\sqrt{N_{s}}$, where $N_{s}$ is the product of the YIG spin density and the intersection of the resonator and YIG film volumes, we infer that the coupling strength does not scale as the square root of the thickness of the latter. In addition, the distribution of the microwave field obtained from the results of electromagnetic simulations, at least for the bare CPW resonator, stretches well above a few microns from the resonator surface [8]. If the magnetic system was paramagnetic or if the mode had uniform precession of the magnetization, all the spins, positioned in a 100-nm-thick layer above the surface of the resonator and located sufficiently near to the central conductor, were uniformly excited by the microwave field. However, the coupling strength estimated using the spin density of the YIG film results much higher than what derived from the experiment; this indicates that in the latter case the precession of the spins is nonuniform, i.e. the spin texture of the hybrid mode is more complex. Thus, while the spin-photon coupling is determined by the vacuum magnetic field of the resonator (Eq. 10), the collective coupling, as suggested by simple calculations, is related to the specific way in which the electromagnetic and spin waves combine in the magnet/superconductor bilayer to shape the hybrid mode.
5 Conclusions
In conclusion, we have studied the microwave response of nonconductive 100-nm-thick YIG films positioned on top of a YBCO superconducting layer. Transmission spectra were acquired at different temperatures either by using a broadband CPW line or a CPW resonator. In the former case, experimental data have been essentially reproduced using well-known formulae and physical quantities describing the spin-wave excitation spectrum of the YIG film. In the latter case, the coupling extracted from the splitting of magnon-photon polaritons amounts to 230 MHz. The evolution of the polaritonic spectrum at different temperatures can be attributed to the effects of the penetration depth of the YBCO superconductor. The comparison with the results obtained with a thicker YIG film suggests a non-trivial role of both system geometry and material parameters in determining the collective coupling strength.
Acknowledgments
This work was partially supported by the project SMILE-SQUIP funded by the Italian NQSTI - National Quantum Science and Technology Institute code PE0000 0023 and by the U.S. Office of Naval Research Award No. N62909-23-1-2079. A.G. acknowledges financial support from PNRR MUR project ECS_00000033_ECOSISTER.